hi i really need someone to solve it fast (2x-1)^2 - (x-4)^2 + 12x - 60 = 0
i know the answer is x = 5/-5 but i need to know how .
thanks !!!

Answer by scott8148(6628)

working the first two terms as the difference of squares

(3x - 5)(x + 3) + 12x - 60 = 0 ___ 3x^2 + 4x - 15 + 12x - 60 = 0 ___ 3x^2 + 16x - 75 = 0

factoring ___ (3x + 25)(x - 3) = 0

the answer that you know, is not correct

you can plug x=5/-5 into the original equation, and see that it does not work